# neurips_2025.bib

### v1

@@ B.2 Timing Experiments
Table 6: Computational cost per inference step for each model across different PDEs, given in milliseconds (ms). Model sizes are also reported for the (Adv/Kdv/SWE) experiments. Each run uses a single NVIDIA RTX 6000 Ada GPU.
@@
The additional overhead of using automatic differentiation, evaluating the operator $\mathcal{J}$, and applying an ODE integrator during inference increases the computational cost of HNFs. We examine this in Table 6 by performing a model rollout and averaging the time per inference step across a batch. Consistent with prior works, FNO models remain a fast neural PDE surrogate. While HNFs may have lower parameter counts, this may be a misleading measure of inference speed. Automatic differentiation calculates the gradient with respect to each input point, therefore the computational cost of this step can scale with the size of the input grid. This can be seen when comparing the Adv and KdV experiments, where the grid size doubles from $n_{x}=128$ to $n_{x}=256$, and the computational cost roughly doubles as well. In 2D experiments, the cost of HNFs is appreciably more than other baselines. We do not anticipate that the finite-difference (FD) schemes used in $\mathcal{J}$ or ODE integrators contribute meaningfully to computational cost, since FD schemes use $\mathcal{O}(M)$ operations, where $M$ is the number of grid points, and numerical integration occurs in constant time.
Appendix C Proof of Linear Functional Approximation
@@
[…]
Let $D\subset\mathbb{R}^{d}$ be a domain. Let $\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$ denote the natural numbers that include zero. Let $\mathcal{X}$ be a Banach space, and $\mathcal{X}^{*}$ be its continuous dual space. Specifically, the dual space contains all continuous, linear functionals $f:\mathcal{X}\rightarrow\mathbb{R}$ with the norm: $||f||_{\mathcal{X}^{*}}=\sup_{x\in\mathcal{X},||x||_{\mathcal{X}}=1}|f(x)|<\infty$.
@@
[…]
@@
[…]
@@
[…]
@@
$||\kappa_{\theta}-\kappa||_{C}\leq\frac{\epsilon}{2M|D|}$
@@
[…]
@@ D.1 Numerical Methods
This can be implemented in constant time by caching prior derivative estimates $\{f(\mathbf{u}^{i}):i<t\}$, allowing the ODE integrator to have a 2nd-order truncation error ($\mathcal{O}(\Delta t^{2})$) while remaining fast. At the first timestep, there are no cached derivatives, therefore a first-order Euler integrator is used: $\mathbf{u}^{t+1}=\mathbf{u}^{t}+\Delta tf(\mathbf{\mathbf{u}}^{t})$. For the given experiments, these integrators are accurate since $\Delta t$ is small enough; to be used with larger timesteps $\Delta t$, higher-order integrators may be needed or smaller, intermediate steps will need to be taken [Zhou_2025]. This adds additional consideration to the training and deployment of derivative-based neural PDE surrogates.
@@
Implementing infinite-dimensional Hamiltonian structure involves approximating the operator $\mathcal{J}$, either through numerical methods or neural operators. We opt for numerical methods, as the operators are simple and in all our tests only involve first-order spatial derivatives $\partial_{x}$. For the 1D Advection equation, we implement a 2nd-order central difference scheme. For the 1D KdV equation, the spatial derivatives exhibit larger gradients, therefore we use an 8th-order central difference scheme with a smoothing stencil to approximate $\partial_{x}$ [snrd]. A similar scheme is implemented in 2D to approximate $\partial_{x}$ and $\partial_{y}$ for the 2D SWE experiments, and additionally, a Savitzky-Golay filter is used to further damp numerical oscillations. These arise since the lack of viscosity creates discontinuities in the shallow-water equations, alternative methods may include flux-limiting or non-oscillatory schemes [HARTEN19973, LIU1994200]. While these implementations work, they still required tuning, as such, the use of a neural operator to approximate $\mathcal{J}$ to make HNFs fully learnable would be an interesting future direction.
Additionally, to calculate temporal derivatives $d\mathbf{u}/dt$ during training, a 4-th order Richardson extrapolation is used to provide accurate derivatives. At the boundaries $t=0,\>t=T$ a one-sided Richardson extrapolation is used [RAHUL200613].
@@
[…]

### v2

@@ B.2 Timing Experiments
Table 6: Computational cost per inference step for each model across different PDEs, given in milliseconds (ms). Model sizes are also reported for the (Adv/Kdv/SWE) experiments. Each run uses a single NVIDIA RTX 6000 Ada GPU. Numerical solvers are run on a AMD Ryzen Threadripper PRO 5975WX 32-Core CPU due to lack of GPU compatibility.
The additional overhead of using automatic differentiation, evaluating the operator $\mathcal{J}$, and applying an ODE integrator during inference increases the computational cost of HNS. We examine this in Table 6 by performing a model rollout and averaging the time per inference step across a batch. Additionally, we report the runtimes of the numerical solvers used. In the case of 1D Advection, the analytical solution is used. Randomly sampled initial conditions can cause variability in the stiffness of the solution, therefore, solver runtimes are averaged across 10 samples. We note that the 2D SWE solver (PyClaw) [31] uses a Fortran compiler and is optimized for hyperbolic PDEs, while the KdV solver is purely Python-based. We note that for a fairer comparison, the numerical solvers should be coarsened to match the accuracy of the neural solvers.
@@
Consistent with prior works, FNO models remain a fast neural PDE solver. While HNS may have a lower parameter count, this may be a misleading measure of inference speed. Automatic differentiation calculates the gradient with respect to each input point, therefore the computational cost of this step can scale with the size of the input grid. This can be seen when comparing the Adv and KdV experiments, where the grid size doubles from $n_{x}=128$ to $n_{x}=256$, and the computational cost roughly doubles as well. In 2D experiments, the cost of HNS is appreciably more than other baselines, although it is still modestly faster than an optimized numerical solver. We do not anticipate that the finite-difference (FD) schemes used in $\mathcal{J}$ or ODE integrators contribute meaningfully to computational cost, since FD schemes use $\mathcal{O}(M)$ operations, where $M$ is the number of grid points, and numerical integration occurs in constant time.
B.3 Ablation Studies
Model Rollout Error
Base 1.248
+ Nonlinear 0.028
+ FiLM 0.016
+ SIREN 0.003
+ Global 0.057
(a) Ablations for 1D Advection
Model Correlation Time
Base 51.75
+ Nonlinear 73.00
+ FiLM 65.25
+ SIREN 73.25
+ Global 151.75
(b) Ablations for 1D KdV
Model Rollout Error
Base 0.113
+ Nonlinear 0.034
+ FiLM 0.030
+ SIREN 0.026
+ Global 0.024
(c) Ablations for 2D SWE-Sines
Table 7: Ablation studies for the HNS architecture. The base architecture is specified by (Linear, Concat, SIREN, Local). Each successive row is cumulative, adding an additional feature to the model. Parameter sizes are approximately constant across ablations.
There are many choices for parameterizing a kernel integral, such as from different kernels (linear, nonlinear), different architectures (MLP, SIREN), different conditioning mechanisms (concat/FiLM) and different receptive fields (local, global). Since this design space is large, we consider the effects of incrementally improving each aspect and report the performance of different HNS models in Table 7. We find that a nonlinear kernel is needed in all cases, since all the Hamiltonians tested are nonlinear. In addition, the use of FiLM conditioning and SIRENs tend to improve performance. Lastly, global conditioning is necessary for the KdV equation, which has nonlocal terms ($\frac{\partial^{2}u}{\partial x^{2}}$) in the Hamiltonian.
B.4 Varying Dataset Sizes
We study the effect of varying the number of samples in the training dataset on HNS performance in Table 8. The training dataset is truncated by a factor of 8, 4, or 2, and evaluated on the same validation dataset. This is repeated across all experiments/PDEs. In general, increasing the dataset size consistently improves model performance. This is more acute for 1D equations. For 2D SWE, the extra spatial dimension may allow for more data per sample relative to the complexity of the dynamics.
Dataset Size Adv KdV SWE-Sines SWE-Pulse
ff=1/8 1.320 71.00 0.0293 0.0243
ff=1/4 0.302 71.25 0.0274 0.0230
ff=1/2 0.075 95.25 0.0254 0.0227
ff=1 0.003 151.75 0.0249 0.0216
Table 8: Rollout error ($\downarrow$) or correlation times ($\uparrow$) of HNS models trained on datasets downsampled by a factor of $f$ (e.g., $f=1/2$ uses half the number of training samples in the training set).
B.5 Discretization Invariance
We experiment with querying models at different discretizations than its training resolution. The results are in Table 9, with the training resolution in bold. Rollout errors are averaged over the validation set, where validation labels are either downsampled or re-solved at a higher resolution using the same initial conditions. Additionally, for 1D Advection, we conducted an experiment in which the model is queried on an unseen grid $x=[0,32],n_{x}=256$, after being trained on a grid $x=[0,16],n_{x}=128$, denoted as $256^{*}$.
(a) Rollout Error at Different Resolutions, 1D Adv
Resolution HNF FNO Unet
64 0.0029 0.064 1.19
128 0.0033 0.064 0.075
256 0.0055 0.050 1.32
256 0.0021 1.34 0.072
(b) Rollout Error at Different Resolutions, 2D SWE (Sines)
Resolution HNF FNO Unet
(64, 64) 0.028 0.066 0.098
(128, 128) 0.029 0.066 0.008
(256, 256) 0.026 0.066 0.110
(c) Rollout Error at Different Resolutions, 2D SWE (Pulse)
Resolution HNF FNO Unet
(64, 64) 0.035 0.173 0.173
(128, 128) 0.024 0.174 0.117
(256, 256) 0.019 0.174 0.225
Table 9: Rollout error of models at different resolutions for 1D Adv and 2D SWE datasets.
In all cases, HNFs are capable of zero-shot super-resolution, as is possible with Neural Operators. Additionally, the error is approximately constant across discretizations and on a new, unseen grid for 1D Advection. Predictably, FNO models have nearly constant error across discretizations. Interestingly, for an extrapolated grid in 1D Advection, FNO struggles to make predictions. For Unet models, the error increases when the grid spacing changes, however for a constant spacing, it is able to extrapolate to unseen grids (i.e. $0,16]\rightarrow[0,32]$ with $dx$ unchanged).
Appendix C Additional Theory
C.1 Proof of Linear Functional Approximation
Interestingly, while development of neural functionals was arrived upon independently, a proof of linear functional approximation exists. In fact, approximating a linear functional with an integral kernel is an intermediate step in operator approximation theorems [77, 78, 79] and its proof comes as a series of lemmas in Kovachki et al. [9]. For the sake of completeness, we provide relevant definitions and cite the necessary lemmas here.
Definitions (Appendix A, Kovachki et al. [9])
@@ C.1 Proof of Linear Functional Approximation
  • $C^{m}(D)=\{f:D\rightarrow\mathbb{R}:\partial^{\alpha}f\in C^{m-|\alpha|_{1}}(D)\>\forall\>0\leq|\alpha|_{1}\leq m\}$
@@
Additionally, we cite definitions of neural networks from Section 8.1, Kovachki et al. [9]. For any $n\in\mathbb{N}$ and $\sigma:\mathbb{R}\rightarrow\mathbb{R}$, a neural network with $n$ layers is given by:
@@
to allow universal approximation. Lastly, functions are restricted to be real-valued, although extensions to vector-valued functions are possible (Section 8.3, Kovachki et al. [9]).
Theorem C.1 (Lemma 30, Kovachki et al. [9]).
@@
$\kappa_{\theta}-\kappa||_{C}\leq\frac{\epsilon}{2M|D|}$
@@
Interestingly, this overlap between neural functionals and neural operators suggests a deeper connection between the two. The integral kernel functional bears many similarities to the integral kernel operator and, in practice, is implemented as an integral kernel operator evaluated at a single output. From an implementation perspective, the two architectures are equivalent in this manner; examining the theoretical basis also suggests that neural functionals are a subset of neural operators when approximating linear functionals. This can be seen by looking at operator approximation theorems (Lemma 22, Kovachki et al. [9]). We refrain from rewriting the theorems here, but focus on the necessary modifications. Specifically, to use a neural operator to approximate a linear functional, the output Banach space $\mathcal{Y}$ is set to $\mathbb{R}$ and intermediate dimensions $J,J^{\prime}$ are set to 1. The linear map $F_{J}:\mathcal{X}\rightarrow\mathbb{R}^{J}$ from the input field $x\in\mathcal{X}$ becomes $F_{J}=w_{1}(x)$, where $w_{1}(x)\in\mathcal{X}^{*}$. Lastly, the finite-dimensional map is defined as $\varphi\in C(\mathbb{R},\mathbb{R})$ and the output map is defined as $G_{J^{\prime}}:\mathbb{R}\rightarrow\mathbb{R}$.
C.2 Connection to the Cylindrical Approximation
Miyagawa and Yokota [56] propose an alternate mechanism to approximate functionals based on the cylindrical approximation. We explore a possible connection with the current method of using an integral kernel.
We consider functionals $F[u]:H\rightarrow\mathbb{R}$, where we restrict functions $u(x)\in H$ to be in a Hilbert space $H$ with inner product $\langle\cdot,\cdot\rangle_{H}$. Since functions can be represented as a sum of basis functions, $u(x)$ can be written as $u(x)=\sum_{k=1}^{\infty}\langle u,\phi_{k}\rangle\phi_{k}(x)$, with an orthonormal basis $\{\phi_{1},\phi_{2},\ldots\}$. Truncating the sum to a finite number of terms $m$ allows the cylindrical approximation of functionals [80, 81, 82]. This is obtained by substituting the truncated series into the functional:
$f(a_{1},a_{2},\ldots,a_{m}):=F[\sum_{k=1}^{m}a_{k}\phi_{k}(x))]$ (15)
where $a_{k}$ are basis coefficients $\langle u,\phi_{k}\rangle$. Using the cylindrical approximation $f$, functional derivatives can be approximated by:
$\frac{\delta F}{\delta u}\approx\sum_{k=1}^{m}\frac{\partial f}{\partial a_{k}}\phi_{k}(x)$ (16)
which converges to the true functional derivative as $m\rightarrow\infty$ [81]. When discretizing the functional derivative, we define a set of $n$ points $\mathbf{x}=[x_{1},x_{2},\ldots,x_{n}]$ at which the basis functions are evaluated, which results in the discrete approximation of the functional derivative:
$\sum_{k=1}^{m}\frac{\partial f}{\partial a_{k}}\phi_{k}(x)\approx\sum_{k=1}^{m}\frac{\partial f}{\partial a_{k}}\phi_{k}(\mathbf{x})=\begin{bmatrix}\sum\frac{\partial f}{\partial a_{k}}\phi_{k}(x_{1})\\ \sum\frac{\partial f}{\partial a_{k}}\phi_{k}(x_{2})\\ \vdots\\ \sum\frac{\partial f}{\partial a_{k}}\phi_{k}(x_{n})\end{bmatrix}$ (17)
We are interested if the gradient of a neural functional that is parameterized by a kernel integral can represent this discrete object, and what conditions are needed to do so. Furthermore, there are two levels of approximation (a truncated basis and a discretized domain) and we are interested if the gradient of a neural functional can recover a functional derivative in its limit. We will work through three cases of increasing complexity, using a linear kernel, a nonlinear kernel, and a nonlinear and global kernel.
Linear Kernel Integrals
A functional parameterized by a linear kernel integral can be written as:
$F_{\theta}[u]=\int_{\Omega}\kappa_{\theta}(x)u(x)\text{d}x\approx\sum_{i=1}^{n}\kappa_{\theta}(x_{i})u_{i}\mu_{i}$ (18)
using a Riemann sum approximation and quadrature weights $\mu_{i}$. Without loss of generality, we can assume some constant quadrature weight $\mu$ and write the gradient of the Riemann sum as:
$\mu\nabla_{u}(\sum_{i=1}^{n}\kappa_{\theta}(x_{i})u_{i})=\mu\begin{bmatrix}\frac{\partial}{\partial u_{1}}(\kappa_{\theta}(x_{1})u_{1})\\ \frac{\partial}{\partial u_{2}}(\kappa_{\theta}(x_{2})u_{2})\\ \vdots\\ \frac{\partial}{\partial u_{n}}(\kappa_{\theta}(x_{n})u_{n}))\end{bmatrix}=\mu\begin{bmatrix}\kappa_{\theta}(x_{1})\\ \kappa_{\theta}(x_{2})\\ \vdots\\ \kappa_{\theta}(x_{n})\end{bmatrix}$ (19)
We can interpret this gradient as the discretization of a single basis function $\kappa_{\theta}=\phi_{1}$, and without the necessary coefficients $\frac{\partial f}{\partial a_{k}}$. Evidently, this is not enough expressivity to approach a sum of $m$ coefficients and basis functions, even as $n\rightarrow\infty$. Intuitively, the output $F_{\theta}[u]$ does not depend on $u$ enough, causing its gradient $\nabla_{u}$ to be too simple. A potential remedy is to modify the kernel to be nonlinear by using $\kappa_{\theta}(x,u(x))$.
Nonlinear Kernel Integrals
Following the same approximation, the gradient of a nonlinear kernel integral can be written as:
$\mu\nabla_{u}(\sum_{i=1}^{n}\kappa_{\theta}(x_{i},u_{i})u_{i})=\mu\begin{bmatrix}\frac{\partial}{\partial u_{1}}(\kappa_{\theta}(x_{1},u_{1})u_{1})\\ \frac{\partial}{\partial u_{2}}(\kappa_{\theta}(x_{2},u_{2})u_{2})\\ \vdots\\ \frac{\partial}{\partial u_{n}}(\kappa_{\theta}(x_{n},u_{n})u_{n}))\end{bmatrix}=\mu\begin{bmatrix}u_{1}\frac{\partial}{\partial u_{1}}\kappa_{\theta}(x_{1},u_{1})\>+\>\kappa_{\theta}(x_{1},u_{1})\\ u_{2}\frac{\partial}{\partial u_{2}}\kappa_{\theta}(x_{2},u_{2})\>+\>\kappa_{\theta}(x_{2},u_{2})\\ \vdots\\ u_{n}\frac{\partial}{\partial u_{n}}\kappa_{\theta}(x_{n},u_{n})\>+\>\kappa_{\theta}(x_{n},u_{n})\end{bmatrix}$ (20)
using the product rule. In general, $\kappa_{\theta}(x_{i},u_{i})$ can be quite complex and nonlinear, however, by construction it only depends on $(x_{i},u_{i})$. We define a function $d_{i}(x_{i})=\frac{\partial}{\partial u_{i}}\kappa_{\theta}(x_{i},u_{i})|_{u_{i}=u_{i}}$, which evaluates the partial derivative at $x_{i}$. This lets us construct at most $n$ basis functions $\{d_{1},d_{2},\ldots,d_{n}\}$, however, this is still not enough expressivity, as each entry of the discretized functional derivative is still restricted to one basis function, albeit a different one at each row. Despite this shortcoming, writing this expansion allows us to see that producing a sum over multiple basis functions requires the kernel to depend globally on $\mathbf{u}=[u_{1},u_{2},\ldots,u_{n}]$, since the gradient of the Riemann sum no longer degenerates to a single term at each row.
Nonlinear Global Kernel Integrals
Using the same approximation, the gradient of a nonlinear, global kernel integral can be written as:
$\mu\nabla_{u}(\sum_{i=1}^{n}\kappa_{\theta}(x_{i},\mathbf{u})u_{i})=\mu\begin{bmatrix}\frac{\partial}{\partial u_{1}}(\sum\kappa_{\theta}(x_{i},\mathbf{u})u_{i})\\ \frac{\partial}{\partial u_{2}}(\sum\kappa_{\theta}(x_{i},\mathbf{u})u_{i})\\ \vdots\\ \frac{\partial}{\partial u_{n}}(\sum\kappa_{\theta}(x_{i},\mathbf{u})u_{i})\end{bmatrix}$ (21)
The expression $\frac{\partial}{\partial u_{j}}(\sum\kappa_{\theta}(x_{i},\mathbf{u})u_{i})$ can be expanded into:
$\frac{\partial}{\partial u_{j}}(\sum_{i=1}^{n}\kappa_{\theta}(x_{i},\mathbf{u})u_{i})$ $=u_{1}\frac{\partial}{\partial u_{j}}\kappa_{\theta}(x_{1},\mathbf{u})\>+\ldots$ (22)
$+\>u_{j}\frac{\partial}{\partial u_{j}}\kappa_{\theta}(x_{j},\mathbf{u})\>+\kappa_{\theta}(x_{j},\mathbf{u})\>+\ldots$
$+\>u_{n}\frac{\partial}{\partial u_{j}}\kappa_{\theta}(x_{n},\mathbf{u})$
In general, we can construct a kernel $\kappa_{\theta}$ such that $\frac{\partial}{\partial u_{j}}\kappa_{\theta}(x_{i},\mathbf{u})$ is different for $i=\{1,\ldots,n\}$. A simple example is provided based on FiLM modulation:
$\kappa_{\theta}(x_{i},\mathbf{u})=[W^{u}\mathbf{u}]_{i}(W^{x}x_{i}+b^{x})+[b^{u}]_{i}$ (23)
where $W^{u}\in\mathbb{R}^{n\times n},W^{x}\in\mathbb{R}^{1x1},b^{u}\in\mathbb{R}^{n},b^{x}\in\mathbb{R}^{1}$ and $\cdot]_{i}$ is an operation that takes the $i$-th row of a column vector. In this case, $\frac{\partial}{\partial u_{j}}\kappa_{\theta}(x_{i},\mathbf{u})$ is equivalent to a function $d_{ij}(x_{i})=W^{q}_{ij}W^{x}x_{i}$. With even deeper and more complex kernels, functions $d_{ij}$ can become more expressive. Without loss of generality, we can also absorb the basis coefficients into $d_{ij}$.
With this formulation, we can see that the $j$-th row of the gradient vector can be written as a sum of $n$ functions $\sum_{i=1}^{n}d_{ij}(x_{i})$, which reduces to the cylindrical approximation when letting $d_{ij}$ be arbitrary. Furthermore, letting $n\rightarrow\infty$ increases the number of functions $d_{ij}$ and coordinates $x_{i}$, and is consistent with the interpretation of simultaneously increasing the number of bases to approach the true functional derivative as well as converting from a discrete to continuous representation.
The cylindrical approximation makes no assumption on the linearity of $F$, therefore, when using a global, nonlinear kernel, a kernel integral can represent a functional derivative. This theoretical insight is also supported by empirical observations. Through ablation studies in Section B.3, we find that modeling complex functional derivatives (such as in the KdV equation) is only possible with a global, nonlinear kernel, and simplifications of the kernel degrade the performance.
C.3 Derivations with Varying Coefficients and Source Terms
Using HNS with PDEs with varying coefficients and source terms is possible on a case-by-case basis, requiring additional effort to derive the Hamiltonian structure. We provide examples below by introducing coefficients to 1D Advection and 1D KdV as well as varying bathymetry (source terms) to 2D SWE.
Varying Coefficients
In the presence of coefficients, the Hamiltonians for 1D Advection and 1D KdV become:
$\mathcal{H}_{adv}[u]=\int_{\Omega}-\frac{c}{2}(u(x))^{2}\text{d}x,\qquad\mathcal{H}_{kdv}[u]=\int_{\Omega}-\frac{\alpha}{6}(u(x))^{3}-\beta u(x)\frac{\partial^{2}u}{\partial x^{2}}(x)\text{d}x,$ (24)
For 1D Advection, one can verify the Hamiltonian structure:
$\frac{\delta\mathcal{H}_{adv}}{\delta u}=-cu(x),\qquad\mathcal{J}(\frac{\delta\mathcal{H}_{adv}}{\delta u})=-c\frac{\partial u}{\partial x}=\frac{\text{d}u}{\text{d}t}$ (25)
A similar calculation recovers the PDE for the 1D KdV equation ($u_{t}+\alpha uu_{x}+\beta u_{xxx}=0$) from Hamilton’s equations by using $\frac{\delta\mathcal{H}_{kdv}}{\delta u}=-\frac{\alpha}{2}u^{2}-\beta u_{xx}$. To accommodate varying coefficients, this suggests training HNS with coefficient information using the loss: $\mathcal{L}(\nabla_{u}\mathcal{H}_{\theta}(\mathbf{u},\mathbf{x},\mathbf{c}),\frac{\delta\mathcal{H}}{\delta u})$. In special cases where the coefficients can be factored out (i.e., $\frac{\delta\mathcal{H}[u,c]}{\delta u}=c\frac{\delta\mathcal{H}[u]}{\delta u}$), once can scale the base model output $\nabla_{u}\mathcal{H}_{\theta}(\mathbf{u},\mathbf{x})$ by $c$ to achieve the same effect.
Varying Source Terms
The 2D shallow-water equations describe free surfaces where the horizontal length scale is significantly larger than the vertical length scale. Within this setup, the free surface can be above a spatially-varying bathymetry $b(x,y)$, or the topography of the bottom surface. Changes in the elevation of the bathymetry can accelerate or decelerate flows, which acts as a source term on the momentum balance [83, 84]. In particular, the shallow-water equations are modified to become:
$\partial_{t}h+\nabla\cdot(\mathbf{v}h)=0,\qquad\partial_{t}\mathbf{v}+\mathbf{v}\cdot\nabla\mathbf{v}=-g\nabla h-gh\nabla b$ (26)
Furthermore, $h(x,y,t)$ now represents the height above the bathymetry $b(x,y,t)$. This source term is represented in the Hamiltonian as:
$\mathcal{H}^{b}[\mathbf{u}]=\mathcal{H}^{b}[v_{x},v_{y},h]=\int_{\Omega}\frac{1}{2}h(v_{x}^{2}+v_{y}^{2})+\frac{1}{2}gh^{2}(1+b)\>\text{d}A$ (27)
and the operator matrix $\mathcal{J}$ is left unchanged (Equation 11). One can check that applying Hamilton’s equations recovers the modified shallow-water equations. When comparing this derivation to one with constant bathymetry, one can show that:
$\frac{\delta\mathcal{H}^{b}[u]}{\delta u}=\frac{\delta\mathcal{H}[u]}{\delta u}+\begin{bmatrix}0\\ 0\\ ghb\end{bmatrix}$ (28)
This suggests that instead of scaling model outputs $\nabla_{u}\mathcal{H}_{\theta}(\mathbf{u},\mathbf{x})$ (as in the case of coefficients), source terms can be potentially captured by adding terms to model outputs. In PDEs where a decomposition like this is not possible, the source term will need to be added to the model input to produce the output $\nabla_{u}\mathcal{H}_{\theta}(\mathbf{u},\mathbf{x},\mathbf{b})$, where bold terms indicate discretized fields/coordinates.
@@ D.1 Numerical Methods
This can be implemented in constant time by caching prior derivative estimates $\{f(\mathbf{u}^{i}):i<t\}$, allowing the ODE integrator to have a 2nd-order truncation error ($\mathcal{O}(\Delta t^{2})$) while remaining fast. At the first timestep, there are no cached derivatives, therefore a first-order Euler integrator is used: $\mathbf{u}^{t+1}=\mathbf{u}^{t}+\Delta tf(\mathbf{\mathbf{u}}^{t})$. For the given experiments, these integrators are accurate since $\Delta t$ is small enough; to be used with larger timesteps $\Delta t$, higher-order integrators may be needed or smaller, intermediate steps will need to be taken [26]. This adds additional consideration to the training and deployment of derivative-based neural PDE solvers.
@@
Implementing infinite-dimensional Hamiltonian structure involves approximating the operator $\mathcal{J}$, either through numerical methods or neural operators. We opt for numerical methods, as the operators are simple and in all our tests only involve first-order spatial derivatives $\partial_{x}$. For the 1D Advection equation, we implement a 2nd-order central difference scheme. For the 1D KdV equation, the spatial derivatives exhibit larger gradients, therefore we use an 8th-order central difference scheme with a smoothing stencil to approximate $\partial_{x}$ [85]. A similar scheme is implemented in 2D to approximate $\partial_{x}$ and $\partial_{y}$ for the 2D SWE experiments, and additionally, a Savitzky-Golay filter is used to further damp numerical oscillations. These arise since the lack of viscosity creates discontinuities in the shallow-water equations, alternative methods may include flux-limiting or non-oscillatory schemes [86, 87]. While these implementations work, they still required tuning, as such, the use of a neural operator to approximate $\mathcal{J}$ to make HNS fully learnable would be an interesting future direction.
Additionally, to calculate temporal derivatives $d\mathbf{u}/dt$ during training, a 4-th order Richardson extrapolation is used to provide accurate derivatives. At the boundaries t=0,t=Tt=0,\>t=T a one-sided Richardson extrapolation is used [88].
@@
The analytical solution to the Advection equation is used to generate its datasets. The numerical solver for the KdV equation is from Brandstetter et al. [24]; due to the stiffness of the PDE, the solutions do not always conserve the Hamiltonian. The WENO schemes used allow for more stable derivatives, but at the cost of numerical viscosity, which can affect the Hamiltonian since it is highly nonlinear. Therefore, we filter the generated KdV solutions and extract 50% of the trajectories with the smallest variation in the Hamiltonian. Similar fluctuations in the Hamiltonian are seen in SWE trajectories generated from PyClaw [31], but the variations are small enough to be ignored.